How many grams of HNO3 are in 75ml of a 8.0M HNO3 solution?


1
Expert's answer
2021-08-25T06:36:40-0400

8.0M means that there are 8 moles in 1L of the solution.

8 moles = 1.0L

x moles = 0.075L


x = "\dfrac{8×0.075}{1.0}" = 0.6 moles


molar mass of HNO3 = "(1+14+(16\u00d73))" = "63\text{ g/mol}"

1 mole of HNO3 = 63 grams

0.6 moles = y grams


y = "\dfrac{63×0.6}1" = 37.8 grams.